1/3

Christian Morgenstern,  Christian Morgenstern,  Christian Morgenstern,  Christian Morgenstern,

Varulven The Banshee (An Approach)

Der Werwolf

The Werewolf

tradukita de Sven Collberg  tradukita de Max Knight  tradukita de Alexander Gross

En varulv lopp från fru m.m. en vacker afton samt begav sig till en bymagisters grav och bad: “Var snäll och konjuge-

ra.

One night, a banshee slunk away from mate and child, and in the

gloom went to a village teacher’s tomb, requesting him: “Inflect me, pray.”

Ein Werwolf eines Nachts entwich von Weib und Kind und sich be-

gab an eines Dorfschullehrers Grab und bat ihn: �Bitte, beuge mich!“ A Werewolf, troubled by his name, Left wife and brood one night and came To a hidden graveyard to enlist The aid of a long-dead philologist.

mig!” Byskolläraren stod upp på blecknamnsskyltens

mässingknopp med svar till den, som satt på

pass med tåligt korslagd tass på tass.

The village teacher climbed up

straight upon his grave stone with its plate and to the apparition said who meekly knelt before the dead:

Der Dorfschulmeister stieg hinauf auf seines Blechschilds Messing-

knauf und sprach zum Wolf, der seine

Pfoten geduldig kreuzte vor dem Toten:

“Oh sage, wake up, please don’t berate

me,” He howled sadly, “Just conjugate me.” The seer arose a bit unsteady Yawned twice, wheezed once, and then was

ready.

“Jag var-ulv 1:a pers. i sing., du, han, hon, den, det var-ulv.

Så vi voro-ulv, I voren. På de voro-ulv kom ingenting.”

“The banSHEE, in the subject’s

place; the banHERS, the possessive case. The banHER, next, is what they call objective case–and that is all.”

�Der Werwolf“ – sprach der gute

Mann, �des Werwolfs, Genitiv sodann, dem Werwolf, Dativ, wie man’s

nennt, den Werwolf, - damit hat’s ein

End!“ “Well, ‘Werewolf’ is your plural past, While ‘Waswolf’ is singularly cast: There’s ‘Amwolf’ too, the present tense, And ‘Iswolf,’ ‘Arewolf’ in this same sense.”

…  …  …  …

2/3

Varulven myste gladåt hur pers. hade följts av sing. och

plur. “Fast”, bad han, “kan det inte ges

ens en enda pers. i tempus presens?”

The banshee marveled at the cases and writhed with pleasure, making

faces, but said: “You did not add, so far, the plural to the singular!”

Dem Werwolf schmeichelten die

Fälle er rollte seine Augenbälle. �Indessen“, bat er, �füge doch zur Einzahl auch die Mehrzahl

noch!“ “I know that–I’m no mental cripple– The future form and participle Are what I crave,” the beast replied. The scholar paused–again he tried:

Bekänna måste vår magister, att därvidlag fanns vissa brister. En varulv sine imperfecto var contradictio in adjecto.

The teacher, though, admitted then that this was not within his ken. “While bans are frequent”, he advi-

sed, “A she cannot be plurized.”

Der Dorfschullehrer aber musste gestehn, dass er von ihr nichts

wusste. Zwar Wölfe gab’s in großer Schar, doch �Wer“ gab’s nur im Sinular.

“A ‘Will-be-wolf?’ It’s just too long:

‘Shall-be-wolf?’ ‘Has-been-wolf?’ Utterly

wrong! Such words are wounds beyond all suture– I’m sorry, but you have no future.”

Varulven satt med tårad lins. Han visste att han fanns och

finns. Därtill kom också fru m.m. som även måste existera.

The banshee, rising clammily, wailed: “What about my family?” Then, being not a learned creature, said humbly “Thanks” and left the te-

acher.

Der Wolf erhob sich tränenblind – er hatte ja doch Weib und Kind! Doch da er keine Gelehrter eben, so schied er dankend und ergeben.

The Werewolf knew better–his sons still

slept At home, and homewards now he crept, Happy, humble, without apology For such folly of philology.

…  …  …  …

3/3

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en la Svedan de Sven Collberg.

Arg-825-1654 (2013-02-13 17:00:43)

Tiu ĉi francigo troviĝas ne la reterjo http: // bertilow. com/ literaturo/ lupfantomo. html# sv .

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en la Anglan de Max Knight (civila nomo: Max Kühnel, ∗1909-06-08 – †1993-08-31).

Arg-825-1660 (2013-02-13 22:36:06)

Tiu ĉi angligo troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html . Pri la tradukinto Max Knight vidu: http: // www. jbeilharz. de/ morgenstern/ morgenstern_poems. html kaj http: // www. 123people. de/ ext/ frm? ti= ersonensuche% 20telefonbuch&search_term= m x% 20k% C3% BChnel&search_ country= DE&st= suc e% 20nach% 20personen&target_url= aHR0cDovL d3dy5saXRlcmF0dXJoYXVzLmF0L2luZGV4LnBocD9pZD02ODc4 .

Verkinto de tiu ĉi Germana poemo estas Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31).

Arg-825-1652 (2013-02-11 18:44:21)

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en la Anglan de Alexander Gross.

Arg-825-1657 (2013-02-13 18:39:45)

Tiu ĉi angligo troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html .